### v1

# On combinatorial invariance of parabolic Kazhdan–Lusztig polynomials

Abstract.
We show that the Combinatorial Invariance Conjecture for Kazhdan–Lusztig polynomials due to Lusztig and to Dyer, its parabolic analog due to Marietti, and a refined parabolic version that we introduce, are equivalent. We use this to give a new proof of Marietti’s conjecture in the case of lower Bruhat intervals and to prove several new cases of the parabolic conjectures.
1991 Mathematics Subject Classification:
05E14, 14M15
GTB was supported by NSF grant DMS-1854512.
CG is supported by a Klarman Postdoctoral Fellowship at Cornell University.
1. Introduction
The Kazhdan–Lusztig polynomial $P_{u,v}(q)\in\mathbb{Z}[q]$, indexed by a pair $u,v$ of elements in a Coxeter group $W$, is a fundamental object in geometric representation theory [12]. These polynomials determine the transition from the standard basis of the Hecke algebra to the Kazhdan–Lusztig basis. When $W$ is the Weyl group of a complex semisimple Lie group $G$, the $P_{u,v}$ give both the Poincaré polynomials of the local intersection cohomology of Schubert varieties in $G/B$ and relate the characters of Verma modules and simple modules of $G$ [2, 9].
The polynomial $P_{u,v}$ is nonzero if and only if $u\leq v$ in Bruhat order; the same is true of the Kazhdan–Lusztig $R$-polynomial $R_{u,v}$, also introduced in [12]. The Combinatorial Invariance Conjecture (Conjecture 1.1) asserts that, remarkably, both families of Kazhdan–Lusztig polynomials are completely determined by the combinatorics of Bruhat order.
Conjecture 1.1 (Lusztig c. 1983; Dyer [11]).
Suppose that for $u_{1},v_{1}\in W_{1}$ and $u_{2},v_{2}\in W_{2}$ the Bruhat intervals $[u_{1},v_{1}]$ and $[u_{2},v_{2}]$ are isomorphic as posets, then:
  • (a)
    $R_{u_{1},v_{1}}=R_{u_{2},v_{2}}$, and
  • (b)
    $P_{u_{1},v_{1}}=P_{u_{2},v_{2}}$.
It is immediate from the definitions (see Definitions 2.2 and 2.3) that the truth of Conjecture 1.1(a) for all pairs of intervals is equivalent to the truth of Conjecture 1.1(b) for all pairs of intervals.
For $J$ a subset of the set $S$ of simple generators of $W$ and $x\in\{-1,q\}$, the parabolic Kazhdan–Lusztig polynomials $P_{u,v}^{J,x}(q)$ generalize the $P_{u,v}$ and, in the case $x=-1$, give the Poincaré polynomials of the local intersection cohomology of Schubert varieties in $G/P_{J}$, where $P_{J}$ is the corresponding parabolic subgroup [10]. Here $u$ and $v$ lie in the parabolic quotient $W^{J}$. The parabolic $R$-polynomials $R_{u,v}^{J,x}(q)$ likewise generalize the $R_{u,v}$. The $P_{u,v}$ and $R_{u,v}$ are the special case $J=\emptyset$.
Let $[u,v]^{J}\coloneqq[u,v]\cap W^{J}$ denote the set of elements from $[u,v]$ lying in $W^{J}$. Brenti, in the case of Hermitian symmetric pairs [6, Corollary 4.8], proved that parabolic Kazhdan–Lusztig polynomials are equal when $[u_{1},v_{1}]^{J}\cong[u_{2},v_{2}]^{J}$ (see also [4, 13]). But examples from [8, 14] show that the naïve extensions of Conjecture 1.1 that this suggests, where one only requires that $[u_{1},v_{1}]^{J_{1}}\cong[u_{2},v_{2}]^{J_{2}}$ or that $[u_{1},v_{1}]\cong[u_{2},v_{2}]$, are false in general. Thus any extension to the parabolic setting must include information about the isomorphism type of the Bruhat intervals as well as information about their intersections with $W^{J}$. Marietti proposed such an extension and proved it for lower intervals.
Conjecture 1.2 (Marietti [14, 15]).
Suppose that for $u_{1},v_{1}\in W_{1}^{J_{1}}$ and $u_{2},v_{2}\in W_{2}^{J_{2}}$ there is a poset isomorphism $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ restricting to an isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$, then:
  • (a)
    $R^{J_{1},x}_{u_{1},v_{1}}=R^{J_{2},x}_{u_{2},v_{2}}$ for $x\in\{-1,q\}$, and
  • (b)
    $P^{J_{1},x}_{u_{1},v_{1}}=P^{J_{2},x}_{u_{2},v_{2}}$ for $x\in\{-1,q\}$.
It was observed by Marietti [15] that the truth of Conjecture 1.2(a) for all pairs of intervals is equivalent to the truth of Conjecture 1.2(b) for all pairs of intervals.
In Conjecture 1.3 below, we propose a refined conjecture in which far less information about $[u,v]^{J}$ is required in order to determine $R^{J,x}_{u,v}$. Let
$A^{J}_{u,v}\coloneqq\{a\in W^{J}\mid u\lessdot a\leq v\}$
denote the set of atoms of $[u,v]$ lying in $W^{J}$.
Conjecture 1.3.
Suppose that for $u_{1},v_{1}\in W_{1}^{J_{1}}$ and $u_{2},v_{2}\in W_{2}^{J_{2}}$ there is a poset isomorphism $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ restricting to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$, then:
  • (a)
    $R^{J_{1},x}_{u_{1},v_{1}}=R^{J_{2},x}_{u_{2},v_{2}}$ for $x\in\{-1,q\}$, and
  • (b)
    $P^{J_{1},q}_{u_{1},v_{1}}=P^{J_{2},q}_{u_{2},v_{2}}$.
Remark 1.4.
The analog of Conjecture 1.3 does not hold for the polynomials $\{P^{J,-1}_{u,v}\}$. For example, letting $W_{1}=W_{2}$ be the symmetric group $S_{4}$, $u_{1}=u_{2}=e$, $v_{1}=s_{1}s_{2}s_{3}$, and $v_{2}=s_{2}s_{1}s_{3}$, both $[u_{1},v_{1}]$ and $[u_{2},v_{2}]$ are isomorphic to the Boolean lattice $B_{3}$. Letting $J_{1}=\{s_{1}\}$ and $J_{2}=\{s_{2}\}$, we have $|A^{J_{1}}_{u_{1},v_{1}}|=|A^{J_{2}}_{u_{2},v_{2}}|=2$. Thus there is an isomorphism $\varphi$ satisfying the hypotheses of the conjecture. However $P^{J_{1},-1}_{u_{1},v_{1}}=1\neq 1+q=P^{J_{2},-1}_{u_{2},v_{2}}$.
Conjecture 1.3(a) implies Conjecture 1.2 since any poset isomorphism $[u_{1},v_{1}]^{J_{1}}\to[u_{2},v_{2}]^{J_{2}}$ in particular restricts to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$. Conjecture 1.2 in turn implies Conjecture 1.1 by taking $J_{1}=J_{2}=\emptyset$. In our first main result, we show that the three conjectures are in fact equivalent. The equivalence of Conjectures 1.1 and 1.2 is already new.
Theorem 1.5.
Conjectures 1.1, 1.2 and 1.3 are equivalent.
In the case of lower intervals, when $u_{1}$ and $u_{2}$ are the identity elements of $W_{1}$ and $W_{2}$ respectively, Conjecture 1.1 was proven by Brenti–Caselli–Marietti [7]. Later, Marietti proved Conjecture 1.2 for lower intervals [14, 15]. We prove Conjecture 1.3 for lower intervals and give a new short proof, relying on Brenti–Caselli–Marietti’s results, of Conjecture 1.2 in this case.
Theorem 1.6.
Conjectures 1.1, 1.2 and 1.3 hold in the case $u_{1}=e_{1}$ and $u_{2}=e_{2}$ are the identity elements of $W_{1}$ and $W_{2}$.
We say a Bruhat interval $[u,v]$ is a short edge interval if all edges $y\to y^{\prime}$ in the Bruhat graph restricted to $[u,v]$ have $\ell(y^{\prime})-\ell(y)=1$. Conjecture 1.1 was proven for short edge intervals by Brenti [5].
Theorem 1.7.
Conjectures 1.1, 1.2 and 1.3 hold when $[u_{1},v_{1}]$ is a short edge interval.
Suppose that $W=S_{n}$ is the symmetric group. We call the Bruhat interval $[u,v]\subset W$ cosimple if $\{\mathbf{e}_{i}-\mathbf{e}_{j}\mid i<j,u\leq v\cdot(ij)\lessdot v\}$ is linearly independent, where the $\mathbf{e}_{i}$ are the standard basis vectors in $\mathbb{R}^{n}$. We call $[u,v]\subset S_{n}$ coelementary if it is isomorphic (as a poset) to some cosimple interval in some symmetric group. The coelementary intervals are related by poset duality to the elementary intervals studied in [1], but the coelementary convention will be more useful for our purposes here.
Theorem 1.8.
Conjecture 1.1(a), Conjecture 1.2(a), and Conjecture 1.3(a) hold when $W_{1}$ and $W_{2}$ are symmetric groups and $[u_{1},v_{1}]$ is a coelementary interval.
Section 2 gives background on Bruhat order and Kazhdan–Lusztig polynomials. In Section 3 we introduce invariant collections of Bruhat intervals and prove Theorem 3.3 which gives general criteria the for combinatorial invariance of parabolic Kazhdan–Lusztig polynomials. Theorems 1.5, 1.6, 1.7 and 1.8 will be shown in Section 4 to follow from Theorem 3.3 and known special cases of Conjecture 1.1.
2. Preliminaries
We refer the reader to [3] for background on Coxeter groups, Bruhat order, and Kazhdan–Lusztig polynomials.
2.1. Parabolic decompositions and Bruhat order
Throughout this work, $W$ will denote a Coxeter group with standard generating set $S$ and length function $\ell$, and $J$ a subset of $S$. We write $W_{J}$ for the parabolic subgroup of $W$ generated by $J$, and $W^{J}$ for the set of minimum-length representatives of the cosets $W/W_{J}$. Each element $w\in W$ may be uniquely decomposed as $w=w^{J}w_{J}$ with $w^{J}\in W^{J}$ and $w_{J}\in W_{J}$. Furthermore, this decomposition satisfies $\ell(w)=\ell(w^{J})+\ell(w_{J})$.
We denote by $\leq$ the (strong) Bruhat order, a partial order on $W$. The following standard fact will be useful (see, e.g. [3, Prop. 2.5.1]).
Proposition 2.1.
For $u,v\in W$ and $J\subseteq S$, if $u\leq v$ then $u^{J}\leq v^{J}$.
We write $[u,v]$ for the closed interval $\{y\in W\mid u\leq y\leq v\}$. The Bruhat graph is the directed graph with vertex set $W$ and directed edges $y\to y^{\prime}$ whenever $y^{\prime}=yt$ for some reflection $t$ and $\ell(y)<\ell(y^{\prime})$.
2.2. Kazhdan–Lusztig polynomials
Definition 2.2 (Kazhdan–Lusztig [12]; Deodhar [10]).
Let $W$ be a Coxeter group, let $J\subset S$, and let $x\in\{-1,q\}$. The family of polynomials $\{R_{u,v}^{J,x}\}_{u,v\in W^{J}}$ is uniquely determined by the following conditions.
  • (i)
    $R_{u,v}^{J,x}=0$ if $u\not\leq v$.
  • (ii)
    $R_{u,u}^{J,x}=1$ for all $u\in W^{J}$.
  • (iii)
    For all $s\in S$ with $\ell(sv)<\ell(v)$ we have:
    $R_{u,v}^{J,x}$ =$=\begin{cases}R_{su,sv}^{J,x},&\text{if $\ell(su)<\ell(u)$}\\ (q-1)R_{u,sv}^{J,x}+qR_{su,sv}^{J,x},&\text{if $\ell(su)>\ell(u)$ and $su\in W^{J}$}\\ (q-1-x)R_{u,sv}^{J,x},&\text{if $\ell(su)>\ell(u)$ but $su\not\in W^{J}$.}\end{cases}$
Definition 2.3 (Kazhdan–Lusztig [12]; Deodhar [10]).
Let $W$ be a Coxeter group, let $J\subset S$, and let $x\in\{-1,q\}$. The family of polynomials $\{P_{u,v}^{J,x}\}_{u,v\in W^{J}}$ is uniquely determined by the following conditions.
  • (i)
    $P_{u,v}^{J,x}=0$ if $u\not\leq v$.
  • (ii)
    $P_{u,u}^{J,x}=1$ for all $u\in W^{J}$.
  • (iii)
    $\deg P_{u,v}^{J,x}\leq\frac{1}{2}(\ell(v)-\ell(u)-1)$ if $u<v$.
  • (iv)
    $q^{\ell(v)-\ell(u)}P^{J,x}_{u,v}(q^{-1})=\sum_{\sigma\in[u,v]^{J}}R^{J,x}_{u,\sigma}P^{J,x}_{\sigma,v}$.
When $J=\emptyset$, it is often omitted from the notation for the $R^{J}$ and $P^{J}$, and in this case the polynomials are independent of the choice of $x\in\{-1,q\}$; these are the ordinary Kazhdan–Lusztig and $R$-polynomials. The following result of Deodhar allows parabolic Kazhdan–Lusztig polynomials to be expressed in terms of ordinary Kazhdan–Lusztig polynomials.
Theorem 2.4.
[10, Prop. 2.12 & Rem. 3.8] For $u,v\in W^{J}$ we have:
  • (a)
    $R^{J,x}_{u,v}=\sum_{w\in W_{J}}(-x)^{\ell(w)}R_{uw,v}$, for $x\in\{-1,q\}$.
  • (b)
    $P^{J,q}_{u,v}=\sum_{w\in W_{J}}(-1)^{\ell(w)}P_{uw,v}$.
3. Invariant collections of intervals
We now define invariant collections of Bruhat intervals. For such collections of intervals we will be able to transfer information about the combinatorial invariance of ordinary $P$- and $R$-polynomials to the parabolic setting (see Theorem 3.3).
Definition 3.1.
Let $\mathcal{I}$ be a collection of Bruhat intervals in Coxeter groups. We say that $\mathcal{I}$ is:
  • upper $R$-invariant (resp. upper $P$-invariant) if for all $[u_{1},v_{1}],[u_{2},v_{2}]\in\mathcal{I}$ and for all poset isomorphisms $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ we have $R_{y,v_{1}}=R_{\varphi(y),v_{2}}$ (resp. $P_{y,v_{1}}=P_{\varphi(y),v_{2}}$) for all $y\in[u_{1},v_{1}]$.
  • fully invariant if for all $[u_{1},v_{1}],[u_{2},v_{2}]\in\mathcal{I}$ and for all poset isomorphisms $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ we have $R_{y,y^{\prime}}=R_{\varphi(y),\varphi(y^{\prime})}$ for all $y,y^{\prime}\in[u_{1},v_{1}]$.
The following fact is immediate from Definition 2.3 (taking $J=\emptyset$ in the definition so that $W^{J}=W$).
Proposition 3.2.
If a collection $\mathcal{I}$ of Bruhat intervals is fully invariant, then for all $[u_{1},v_{1}],[u_{2},v_{2}]\in\mathcal{I}$ and for all poset isomorphisms $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ we have $P_{y,y^{\prime}}=P_{\varphi(y),\varphi(y^{\prime})}$ for all $y,y^{\prime}\in[u_{1},v_{1}]$.
Proposition 3.2 implies that fully invariant collections are in particular upper $R$-invariant and upper $P$-invariant.
Theorem 3.3 below is our most general result. Theorems 1.5, 1.6, 1.7 and 1.8 will all be shown to follow from it.
Theorem 3.3.
Let $\mathcal{I}$ be a collection of Bruhat intervals in Coxeter groups. Let $[u_{1},v_{1}],[u_{2},v_{2}]\in\mathcal{I}$ with $u_{1},v_{1}\in W_{1}^{J_{1}}$ and $u_{2},v_{2}\in W_{2}^{J_{2}}$ and let $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ be a poset isomorphism restricting to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$.
  • (a)
    If $\mathcal{I}$ is upper $R$-invariant, then $R^{J_{1},x}_{u_{1},v_{1}}=R^{J_{2},x}_{u_{2},v_{2}}$ for $x\in\{-1,q\}$.
  • (b)
    If $\mathcal{I}$ is upper $P$-invariant, then $P^{J_{1},q}_{u_{1},v_{1}}=P^{J_{2},q}_{u_{2},v_{2}}$.
If moreover $\varphi$ sends $[u_{1},v_{1}]^{J_{1}}$ to $[u_{2},v_{2}]^{J_{2}}$, then:
  • (c)
    If $\mathcal{I}$ is fully invariant, we have $P^{J_{1},x}_{u_{1},v_{1}}=P^{J_{2},x}_{u_{2},v_{2}}$ and $R^{J_{1},x}_{u_{1},v_{1}}=R^{J_{2},x}_{u_{2},v_{2}}$ for $x\in\{-1,q\}$.
We first prove a key lemma.
Lemma 3.4.
Let $u,v\in W^{J}$ with $u\leq v$. Then
(1) $uW_{J}\cap[e,v]=\{y\in[u,v]\mid y\not\geq a\text{ for all }a\in A^{J}_{u,v}\},$
where $e$ is the identity element of $W$.
Proof.
Suppose $y=uw\in uW_{J}\cap[e,v]$. Since $u\in W^{J}$, we have $u<uw$, so $y\in[u,v]$. If we had $y\geq a$ for some $a\in A^{J}_{u,v}$, then we would have $y^{J}\geq a^{J}=a>u$ by Proposition 2.1, but this contradicts the uniqueness of the parabolic decomposition $y=uw$. Thus $y$ belongs to the right hand side of (1).
Suppose now that $y\in[u,v]$ satisfies $y\not\geq a$ for all $a\in A^{J}_{u,v}$; we must show that $y\in uW_{J}$. By Proposition 2.1, we have $y^{J}\in[u,v]^{J}$. Since $[u,v]^{J}$ is graded by length [3, Cor. 2.5.6] and has unique minimal element $u$, if $y^{J}>u$ then $y^{J}\geq a$ for some $a\in A^{J}_{u,v}$. But this cannot be the case, since we would have $a\not\leq y\geq y^{J}\geq a$. Thus we must have $y^{J}=u$, and so $y=uy_{J}\in uW_{J}$. ∎
Proof of Theorem 3.3.
Let $[u_{1},v_{1}],[u_{2},v_{2}]\in\mathcal{I}$ with $u_{1},v_{1}\in W_{1}^{J_{1}}$ and $u_{2},v_{2}\in W_{2}^{J_{2}}$ and let $\varphi:[u_{1},v_{1}]\to[u_{2},v_{2}]$ be a poset isomorphism restricting to a bijection $A^{J_{1}}_{u_{1},v_{1}}\to A^{J_{2}}_{u_{2},v_{2}}$. The key observation is that the right-hand side of (1) from Lemma 3.4 is preserved by $\varphi$ (a fact which is not clear a priori for the left-hand side). Indeed, we have
$\varphi\left(\{y_{1}\in[u_{1},v_{1}]\mid\forall a\in A^{J_{1}}_{u_{1},v_{1}},~{}y_{1}\not\geq a\}\right)$
=$=\{\varphi(y_{1})\in[u_{2},v_{2}]\mid\forall a\in\varphi(A^{J_{1}}_{u_{1},v_{1}}),~{}\varphi(y_{1})\not\geq a\}$
(2) =$=\{y_{2}\in[u_{2},v_{2}]\mid\forall a\in A^{J_{2}}_{u_{2},v_{2}},~{}y_{2}\not\geq a\}.$
Where in the first equality we have used that $\varphi$ is a poset isomorphism and in the second equality we have used the fact that $\varphi$ sends $A^{J_{1}}_{u_{1},v_{1}}$ to $A^{J_{2}}_{u_{2},v_{2}}$.
  • (a)
    Suppose that $\mathcal{I}$ is upper $R$-invariant. By Theorem 2.4(a) we have
    $R^{J_{1},x}_{u_{1},v_{1}}=\sum_{w\in W_{J_{1}}}(-x)^{\ell(w)}R_{u_{1}w,v_{1}}.$
    Since $R_{u_{1}w,v_{1}}=0$ unless $u_{1}w\leq v_{1}$, this sum is the same as
    $\sum_{y_{1}\in u_{1}W_{J_{1}}\cap[e,v_{1}]}(-x)^{\ell(y_{1})-\ell(u_{1})}R_{y_{1},v_{1}}=\sum_{\begin{subarray}{c}y_{1}\in[u_{1},v_{1}]\\ \forall a\in A^{J_{1}}_{u_{1},v_{1}},~{}y_{1}\not\geq a\end{subarray}}(-x)^{\ell(y_{1})-\ell(u_{1})}R_{y_{1},v_{1}}.$
    We have $R_{y_{1},v}=R_{\varphi(y_{1}),v_{2}}$ by upper $R$-invariance and hence by (2) we conclude
    $R^{J_{1},x}_{u_{1},v_{1}}=\sum_{\begin{subarray}{c}y_{2}\in[u_{2},v_{2}]\\ \forall a\in A^{J_{2}}_{u_{2},v_{2}},~{}y_{2}\not\geq a\end{subarray}}(-x)^{\ell(y_{2})-\ell(u_{2})}R_{y_{2},v_{2}}=R^{J_{2},x}_{u_{2},v_{2}}.$
  • (b)
    If instead $\mathcal{I}$ is upper $P$-invariant, then we can apply Theorem 2.4(b) and argue as above to conclude that $P^{J_{1},q}_{u_{1},v_{1}}=P^{J_{2},q}_{u_{2},v_{2}}$.
  • (c)
    Suppose now that $\varphi$ sends $[u_{1},v_{1}]^{J_{1}}$ to $[u_{2},v_{2}]^{J_{2}}$ and that $\mathcal{I}$ is fully invariant. For any $u_{1}^{\prime},v_{1}^{\prime}\in[u_{1},v_{1}]^{J_{1}}$ the restriction $\varphi|_{[u_{1}^{\prime},v_{1}^{\prime}]}$ is a poset isomorphism onto its image $[u_{2}^{\prime},v_{2}^{\prime}]$, where $u_{2}^{\prime}\coloneqq\varphi(u_{1}^{\prime}),v_{2}^{\prime}\coloneqq\varphi(v_{1}^{\prime})$. Furthermore, it sends $A^{J_{1}}_{u_{1}^{\prime},v_{1}^{\prime}}$ to $A^{J_{2}}_{u_{2}^{\prime},v_{2}^{\prime}}$. Thus, by the arguments in part (a), we have that
    (3) $R^{J_{1},x}_{u_{1}^{\prime},v_{1}^{\prime}}=R^{J_{2},x}_{u_{2}^{\prime},v_{2}^{\prime}}.$
    If $u_{1}=v_{1}$ then $u_{2}=v_{2}$ and we have $P^{J_{1},x}_{u_{1},v_{1}}=P^{J_{2},x}_{u_{2},v_{2}}=1$. If $u_{1}<v_{1}$ then
    (6) $q^{\ell(v_{1})-\ell(u_{1})}P^{J_{1},x}_{u_{1},v_{1}}(q^{-1})-P^{J_{1},x}_{u_{1},v_{1}}(q)$ $=\sum_{\begin{subarray}{c}\sigma_{1}\in W_{1}^{J_{1}}\\ u_{1}<\sigma_{1}\leq v_{1}\end{subarray}}R^{J_{1},x}_{u_{1},\sigma_{1}}(q)P^{J_{1},x}_{\sigma_{1},v_{1}}(q)$
    (9) $=\sum_{\begin{subarray}{c}\sigma_{2}\in W_{2}^{J_{2}}\\ u_{2}<\sigma_{2}\leq v_{2}\end{subarray}}R^{J_{2},x}_{u_{2},\sigma_{2}}(q)P^{J_{2},x}_{\sigma_{2},v_{2}}(q)$
    (10) $=q^{\ell(v_{2})-\ell(u_{2})}P^{J_{2},x}_{u_{2},v_{2}}(q^{-1})-P^{J_{2},x}_{u_{2},v_{2}}(q).$
    Here we have used (3) and have assumed by induction on the height of the intervals that $P^{J_{1},x}_{\sigma_{1},v_{1}}=P^{J_{2},x}_{\varphi(\sigma_{1}),v_{2}}$ for $u_{1}<\sigma_{1}\leq v_{1}$. The fact that $\varphi$ is a poset isomorphism implies that $\ell(v_{1})-\ell(u_{1})=\ell(v_{2})-\ell(u_{2})$. Together with the degree bound from Definition 2.3(iii), equation (10) then implies that $P^{J_{1},x}_{u_{1},v_{1}}=P^{J_{2},x}_{u_{2},v_{2}}$.
4. Applications of Theorem 3.3
The proofs of the remaining theorems follow from Theorem 3.3.
Proof of Theorem 1.5.
It was noted in the introduction that Conjecture 1.3(a) implies Conjecture 1.2 which in turn implies Conjecture 1.1. Thus it suffices to show that Conjecture 1.1 implies Conjecture 1.3.
Suppose that Conjecture 1.1 holds. This implies that for all $W_{1}$ and $W_{2}$ the collection $\mathcal{I}$ of all Bruhat intervals in $W_{1}$ and $W_{2}$ is fully invariant and therefore is, in particular, upper $R$-invariant and upper $P$-invariant. By Theorem 3.3(a) and (b), we have Conjecture 1.3 (a) and (b) respectively. ∎
Proof of Theorem 1.6.
It follows from [7, Thm. 7.8] that the collection of lower intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.3 (a) and (b) for lower intervals, and applying Theorem 3.3(c) yields this case of Conjecture 1.2. ∎
Proof of Theorem 1.7.
It follows from [5, Thm. 6.3] that the collection $\mathcal{I}$ of short edge intervals in $W_{1}$ and $W_{2}$ is fully invariant. Applying Theorem 3.3(a) and (b) proves Conjecture 1.3 (a) and (b) for intervals from $\mathcal{I}$, and applying Theorem 3.3(c) yields this case of Conjecture 1.2. ∎
Proof of Theorem 1.8.
It was shown in [1, Thm. 1.6] that Conjecture 1.1(a) holds when $W_{1}$ and $W_{2}$ are symmetric groups and $[u_{1},v_{1}]$ and $[u_{2},v_{2}]$ are elementary intervals. Multiplication by the longest element $w_{0}$ of $S_{n}$ induces an antiautomorphism of Bruhat order and sends elementary intervals to coelementary intervals. Since we also have $R_{u,v}=R_{w_{0}v,w_{0}u}$ for all $u\leq v\in S_{n}$ [3, Exer. 5.10(b)], the result of [1] also implies that Conjecture 1.1(a) holds for coelementary intervals in symmetric groups. Upper subintervals $[y,v]$ of coelementary intervals $[u,v]$ are easily seen to be coelementary themselves. Thus the collection $\mathcal{I}$ of coelementary intervals in symmetric groups is upper $R$-invariant. Applying Theorem 3.3(a) yields Conjecture 1.2(a) and Conjecture 1.3(a). ∎
Acknowledgements
We are grateful to Mario Marietti for his helpful comments.